### v1

# JWST imaging of the Pleiades: anisotropy of turbulence in the cold neutral medium

JWST imaging of the Pleiades: anisotropy of turbulence in the cold neutral medium
Interstellar medium studies rely on magnetohydrodynamic (MHD) turbulence as a framework for interpretation. In this context, the statistical characterization of interstellar observations is of prime importance. We open a new perspective on diffuse interstellar matter by analyzing James Webb Space Telescope (JWST) observations of the Pleiades nebula with NIRCam. These observations are remarkable in that they provide a microscope view at the cold neutral medium (CNM) with a spatial resolution of 0.2 mpc (40 au). A two-dimensional Fourier analysis is used to characterize the structure of PAH emission in regions near and far from the Pleiades star Merope. To produce maps of the interstellar emission, stars and galaxies are filtered out. The final step in the data cleaning involves subtracting a component, in Fourier space, which we infer to be a residual of the near-infrared cosmic background. The PAH emission power spectra are highly anisotropic. They are well fitted with a break-free power-law, suggesting that we do not observe a specific scale for energy dissipation. Power-law indices are -3.5 near Merope and -3 in the more distant field. The magnetic field orientation, as derived from the Planck dust polarization data, aligns with the PAH anisotropy. The power anisotropy is constant across scales. These findings are discussed in relation to interstellar turbulence that may be driven by the Pleiades stars. The JWST observations of the Pleiades offer a new viewpoint for comparing observations and theoretical models, as they examine physical scales at which turbulence in the CNM is subsonic and decoupled from the thermal instability. The observations may indicate that the turbulent energy cascade in the CNM is anisotropic.
Key Words.:
ISM, Magnetic fields, Turbulence1 Introduction
Studies of the cold interstellar medium (ISM) frequently rely on numerical simulations of magnetohydrodynamic (MHD) turbulence as a theoretical model (Hennebelle and Falgarone, 2012; Burkhart, 2021). In this context, the statistical analysis of interstellar observations is crucial to interpret observations (Richard et al., 2025). This paper opens a new perspective, using observations of the Pleiades nebula by the James Webb Space Telescope (JWST) with the Near-Infrared Camera (NIRCam).
The Pleiades reflection nebula is an iconic astronomical object familiar to a wide audience and also a prime scientific target to study the dusty interstellar medium (Gibson and Nordsieck, 2003a, b; Gibson, 2007). We are observing a chance encounter between a star cluster and diffuse molecular gas (Gordon and Arny, 1984; Herbig and Simon, 2001; Ritchey et al., 2006). The brightest members of the Pleiades cluster are late B-stars at a distance of 135 pc (Abramson, 2018), illuminating their surrounding ISM over a few parsecs (Gibson and Nordsieck, 2003a). The stellar illumination enables the ISM structure to be observed at smaller angular scales than is usually possible. The reflection nebulosity offers a spectacular view at the filamentary structure of the diffuse ISM.
Many studies have relied on dust continuum observations to determine the power spectrum of interstellar matter (Gautier et al., 1992; Ingalls et al., 2004; Miville-Deschênes et al., 2007, 2010, 2016). The advent of Planck made it possible to extend this type of analysis to polarized dust emission (Planck Collaboration, 2020a), which established the statistical alignment of diffuse interstellar structures with the magnetic field direction on parsec scales (Clark et al., 2015; Planck Collaboration, 2016c). This result has been interpreted as a hallmark of magnetized interstellar turbulence (Caldwell et al., 2017; Kandel et al., 2017). In a broader context, the James Webb Space Telescope (JWST) offers a unique opportunity. Its angular resolution and sensitivity to the emission of polycyclic aromatic hydrocarbons (PAHs) enable imaging of the cold interstellar medium (ISM) at finer scales than was previously possible. At the distance of the Pleiades, the angular resolution of NIRCam corresponds to a spatial scale of 0.2 mpc, at which we may be able to identify signatures of dissipation of turbulence, in particular through ion-neutral friction (Burkhart et al., 2015; Hu et al., 2024). Although there is observational evidence of structure in the cold neutral medium (CNM) at these scales, existing observations are too sparse to conduct a statistical study (Heiles, 1997; Stanimirović and Zweibel, 2018).
The purpose of this work is to demonstrate and capitalize on the potential of JWST for diffuse ISM studies. We use PAH emission to investigate the structure of the cold magnetized ISM. In fact, PAHs are tightly coupled to the gas through collisions (Omont, 1986) and to magnetic field lines. For physical conditions in the Pleiades nebula – specifically the gas density and temperature, and the radiation field strength – PAHs exist partly as cations and partly in neutral form (Flagey et al., 2006). However, both charged and neutral PAHs are coupled to magnetic field lines because their charge state varies on timescales much shorter than their drag time through collisions with the gas (Lepp and Dalgarno, 1988).
The paper is organized as follows. Section 2 introduces the JWST observations and outlines the data processing steps to generate images of diffuse emission. Our two-dimensional Fourier analysis of these images is described in Sect. 3 and Sect. 4. In Sect. 5, we discuss the results of the data analysis in relation to interstellar turbulence. Finally, Section 6 summarizes our main results. The paper includes three appendices. Appendix A details and illustrates the methods used to remove stars and mask galaxies from the JWST images. Appendix B provides a broader context for the JWST observations, in which we characterize the magnetized interstellar matter associated with the Pleiades nebula using Planck data. In Appendix C, we use the observations to characterize the physical conditions and demonstrate that the JWST observations of the Pleiades probe the structure of the cold neutral medium.
2 Imaging the diffuse emission from PAHs
This section introduces the JWST observations (Sect. 2.1) and outlines the data processing steps that were performed to derive maps of diffuse emission from PAHs (Sect. 2.2).
2.1 JWST Observations
The NIRCam instrument (Rieke et al., 2023) was used to image two fields, each measuring by as part of the GO 2143 program. Field 1 is centered approximately 2′ to the East of the bright Pleiades star Merope (23 Tau), and Field 2 about 10′ to the South-East of Merope. The top images in Fig. 1 show the location of the two fields within the Pleiades. The two images at the bottom were produced by the JWST pipeline. For this paper, we used version 1.18.0 of the calibration software.
The two fields were imaged in the F212N filter, centered on the 1-0 S(1) H2 line, and the F335M filter that includes the 3.3-m emission bands from polycyclic aromatic hydrocarbons (PAHs). We also obtained observations at both sky positions in the F200W and F277W filters. The observations in the short- and long wavelength channels are executed simultaneously, in the spectral filters F212N with F335M and F200W with F277W. For each pointing, the integration time was for the H2 and PAH observations, and for the wide filters. In this study, we only analyze the F335M data. The data obtained using the three additional filters will be presented in a future paper.
The long channel of NIRCam comprises two square detectors, each with 2000 pixels square and a field of view of . We used the FULLBOX 6TIGHT111NIRCam page at https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-observing-modes/ dithering pattern to obtain multiple offset pointings of the camera and cover the rectangular field without gaps. The dithering also mitigates bad pixels and flat-field uncertainties; it significantly reduces the data noise on the smallest angular scales, where brightness variations are the smallest. We did not use secondary subpixel dithers because the fine sampling of the point-spread function is not essential to our observations.
2.2 Maps of diffuse emission
In order to study the PAH emission, we need to separate the diffuse emission in the F335M filter image from numerous point sources (mostly stars) and compact sources (mostly galaxies). We proceed in four main steps, which we outline here. More details are given in Appendix A.
We model and subtract the large diffraction spikes from two bright stars Merope and PQ Tau. These two stars are outside the JWST fields, but their spikes are prominent structures in Field 1. One of the Merope spikes is also visible in Field 2.
We subtract from a first approximation the diffuse emission using a multiscale median filter. This step yields background-subtracted images, which we then use to identify and remove point sources and compact sources located within the JWST fields.
We subtract point sources using an empirical template of the JWST point spread function (PSF). This template was obtained by stacking background-subtracted images centered on point sources to subpixel accuracy. Point sources are subtracted by fitting the PSF template over small images centered on each source. We do not use the core of the point sources in this fit because it is often saturated.
Compact extended sources — including galaxies and the saturated cores of bright stars — are identified on maps after background, spike, and point source subtraction, and are then masked. The sky area we mask is set by a brightness threshold applied to maps in which spikes, multi-scale background, and point sources have been subtracted (see Appendix A). The masked areas are filled in with the multi-scale background maps.
To assess systematic uncertainties associated with the data cleaning, we check how sensitive the results of our data analysis are to the brightness threshold used to create the mask. For the fiducial mask, 9% of the sky pixels are masked. We use two alternative masks, the light mask for which this fraction is 5% and the heavy mask for which it is 14%.
For statistical analysis, we divide the dither pointings into two sets and , which we use to produce a pair of images for each of the two fields. The two images in each pair cover the same sky region, but have independent statistical noise. Each of the four images (2 pairs) is processed separately by following steps (1) to (4). Separate background maps and masks are computed for each image. After step (4), the difference between the two images in each pair allows us to identify and mask a small number of glitches present in one of the two images. Mathematically, for each field, we obtain two independent maps of diffuse emission and from sky images and with spikes and stars subtracted:
where and denote masks and and the background maps.
For power spectrum analysis, we cut out two square images, with 2000 pixels on each side, in each field. The positions of these cuts are drawn with solid black lines in the JWST images displayed in Fig. 1. The center positions and orientation of these images can be found in Table 1. We analyze eight square images (four pairs) from which diffraction spikes, stars, galaxies, and glitches have been removed. The maps of diffuse emission in each of the four pairs are presented in Fig. 2. The four images show a similar pattern with nearly parallel striations. The JWST/NIRCam pixel size corresponds to mpc (8.5 au) at the distance of the Pleiades. The full size of each square image is 80 mpc. Compared with other observational means, one may say that the JWST offers us a microscope view at the diffuse ISM.
Pairs of square images are used to compute cross-power spectra unbiased by data noise. To emphasize and demonstrate the importance of data cleaning, we calculated the cross-power spectra at each stage for Field 1 Cut 1 as an example. The results of this analysis are shown in Fig. 3. The top spectrum represents the unprocessed JWST image that features spikes, stars, and galaxies. In contrast, the bottom spectrum corresponds to the final image of the diffuse emission. The intermediate curves indicate the cross-power spectra of the maps of spikes, stars, and galaxies that were subtracted from the original data. Data cleaning leads to a decrease in the amplitude of the spectra by up to 4 orders of magnitude for the highest .
The mean brightness has been subtracted from the JWST images because it is a model value (Rigby et al., 2023) introduced in the data pipeline that does not separate interstellar emission from the zodiacal, stellar, and extragalactic backgrounds. To estimate the mean PAH emission, we measured the mean brightness averaged over each image field of view at m, using the Wide-Field Infrared Explorer (WISE) data (Wright et al., 2010). The values listed in Table 1 represent the emission from the Pleiades nebula with the large-scale Galactic contribution, measured off the nebula, subtracted out. The brightness ratios between the diffuse Galactic emission in the F335M NIRCam filter and the WISE m filter, values listed in Table 1, were measured by correlating the structure of the 12m emission with the multiscale background F335M maps. These values are consistent with dust emission models (Compiègne et al., 2011; Draine and Li, 2007) and observations222The color ratio is related to the size distribution of PAHs. The diffuse ISM values, derived from the analysis of the DIRBE and Spitzer/IRAC data, are in the range 2 to (Dwek et al., 1997; Flagey et al., 2006).
| Image | RA | Dec | PA | ||
| J2000 | J2000 | deg | MJy sr-1 | ||
| (a) | (a) | (b) | (c) | (d) | |
| Field1 | -15.4 | ||||
| Cut 1 | 3:46:26.61 | 23:58:18.5 | |||
| Cut 2 | 3:46:29.55 | 23:55:27.4 | |||
| Field2 | -16.1 | ||||
| Cut 1 | 3:46:49.64 | 23:51:56.3 | |||
| Cut 2 | 3:46:52.72 | 23:49:5.7 |
Notes. (a) Coordinates of image centers. (b) Position angle of the long axis of the JWST mosaic with respect to the North direction measured positive towards East (counterclockwise direction). (c) Mean emission at m averaged over each of the Pleiades images, measured using the WISE full resolution image. The background contribution is subtracted. (d) Emission ratio derived from a linear correlation of the NIRCam F335M background image and the WISE data.
3 Fourier analysis
This section presents the Fourier analysis of the images of diffuse emission shown in Fig. 2.
3.1 Computation and statistical uncertainties
We compute the two-dimensional cross-power spectrum for a pair of square images and introduced in Eq.LABEL:eq:im_painting.
where the wavenumber and the angle , measured from the Celestial North modulo , are polar coordinates in the Fourier plane. denotes the real part, the hat symbol the Fourier transform, and the asterisk the complex conjugate. represents the image apodization applied to avoid edge effects in the Fourier transform333We use the Tukey function in the SciPy python package with its default parameters..
We denote the mean value of , within bins of width and , and the cross-power spectrum averaged over angles. To estimate uncertainties on and , we use the following formulae :
where is the noise power spectrum of images and , is the number of Fourier modes in the bin , and is the effective area of the sky covered by the maps after apodization. We refer to Knox (1995) and Hivon et al. (2002) for a mathematical derivation and the numerical assessment of these formulae, and to Viero et al. (2013) and Córdova Rosado et al. (2024) for their application to the analysis of infrared sky maps. To apply this formula, we require the noise power spectra for each pair of square maps. These were estimated from the power spectrum of the difference , for Fields 1 and 2 separately. The factor 2 arising from the difference was taken into account.
The left term in Eq. 3 represents the cosmic variance, the minimum uncertainty on the sky spectrum estimate for noise-free data, the right term is the variance of the data noise, and the middle term the random correlation between sky emission and data noise. The left term dominates for the lowest wavenumbers, where the maps comprise a small number of independent modes, and the right term dominates for the highest wavenumbers, where the signal-to-noise ratio on the sky emission is the lowest.
Figure 4 shows the cross power spectra of the four JWST images. Hereafter, the wavenumber expressed in units of in the bottom axis is converted into physical scales in au on the top axis for a distance to the Pleiades of 135 pc (Melis et al., 2014; Abramson, 2018). The spectra are corrected for the signal attenuation by the telescope beam, using the power spectrum of the point spread function (PSF) image produced with the STPSF Python package for our NIRCam observations in the F335M filter. The PSF varies across the field of the camera, but this effect is ignored because it is smeared out by the raster mapping of the sky. The spectra in Figure 4 all show a flattening at high . This is not due to data noise, because the noise is not correlated between maps and .
3.2 Anisotropy model
In order to analyze the anisotropy of the diffuse emission, we average into logarithmic bins of k with and linear bins of , modulo , with , which do not overlap. We denote and the bin centers and the binned value of . The uncertainty on is estimated using Eq. 3. We model the angular dependence of with a Gaussian profile. Mathematically our data model is as follows:
The angle defines the orientation, perpendicular to the striations seen in the images, for which the power is maximum, and measures the dispersion of this orientation over the images. represents the ratio of maximum to minimum power while the minimum power within the bin. We perform a global fit of with values of and independent of . These two model parameters are listed in Table 2 for each of the four images in Fig. 2. The values of match the dust polarization angles measured at the image centers using Planck polarization data at 353 GHz (see Appendix B and Fig. 21). In emission, the dust polarization angle is perpendicular to the projection of the magnetic field on the plane of the sky. As measures the orientation perpendicular to the striations, we conclude that the filamentary structure of the PAH emission is aligned with the magnetic field projection. This is similar to what is observed on parsec scales with Planck data (Clark et al., 2015; Planck Collaboration, 2016c).
Figure 5 presents the angular profiles of the power spectrum amplitude obtained for Cut 1 in Field 1. The profiles are shown for five -bins, equally spaced in log() from large to small angular scales from top to bottom; the dashed lines represent the model fit to the data. We note that the profile appears to be less anisotropic for high wavenumbers444The fit shown in Fig. 5 allows for small variations of and between the five bins..
The model allows us to derive the power spectra for the orientations parallel and perpendicular to the striations in the sky image:
where is the power spectrum of the JWST PSF. The division by corrects the spectra for the attenuation of the beam. The data error bars are propagated to and by the fitting procedure. Our definition of corresponds to the power spectrum of the emission parallel to the mean orientation of the striations555This is true provided that the Gaussian in Eq. 5 is negligible at of , a condition that is satisfied for the values of in Table 2.. As Cho and Vishniac (2000) pointed out in their analysis of MHD turbulence simulations, the Fourier transformation only partially measures the three-dimensional anisotropy of the PAH emission because the striations orientation varies across the sky and along the line of sight. The angular widths of the profiles in Fig. 5 reflect the dispersion of the orientation of the striations across the image. These widths are significant. They exceed the dispersion of the polarization angles () reported by Planck Collaboration (2016a) in their analysis of Planck data in the southern Galactic cap.
3.3 CIB residual
Figures 4 and 5 indicate that the power spectra include an isotropic emission component, which prevails at high values. We characterize this component, which we identify as an extragalactic residual left after data cleaning.
| Mask | |||
| [Jy2 sr-1] | |||
| (a) | (b) | (c) | |
| Fiducial | 2.8 | ||
| Light | 24 | ||
| Heavy | 13 |
Notes. (a): Amplitude of power-law at wavenumber pix. (b): Spectral index of the power-law. (c): Reduced of the fit.
The isotropic component is most readily apparent in the spectra of Field 2 Cuts 1 and 2 displayed in Fig. 6, because in these spectra the PAH contribution is the smallest. One can clearly see that the two spectra flatten toward high . Figure 6 shows fits of the spectra with a model including two power laws: a steep one representing the PAH contribution and a shallow one representing the isotropic component. For the isotropic component, the same power law is used for Cuts 1 and 2. We performed the data fit for the fiducial, light and heavy masks. The amplitudes and spectral indices of the isotropic component are listed in Table 3. The amplitude and spectral index of the isotropic component are consistent for the fiducial and heavy masks. The reduced of the fit is good for the fiducial mask but poor for the light and heavy masks. For the light mask, the isotropic component is dominant for all and the fit does not succeed in separating the PAH and the isotropic components. For the heavy mask, the amplitude of the isotropic component differs significantly between Cuts 1 and 2.
The spectral index of the isotropic component, about , coincides with the value measured for the clustering of faint extragalactic sources combined into a diffuse extragalactic background. The cosmic near-infrared background (CIB) was detected with IRAC Spitzer data. The measurements reported by Kashlinsky et al. (2012) for the IRAC m channel can be directly compared with the results of our JWST analysis. The spectral index of the isotropic component is consistent with the Spitzer CIB power spectra, and the amplitude is one order of magnitude smaller. We conclude that the isotropic component is a residual CIB, which remains after the masking of extragalactic sources666We note that the spectra in Fig. 6 do not exhibit flattening at high , which could be interpreted as shot noise from faint sources at the JWST detection limit..
| Image | Mask | |||
| [Jy2 sr-1] | [MJy sr-1] | |||
| (a) | (b) | (c) | (d) | |
| Cut 1 | F | |||
| L | ||||
| H | ||||
| Cut 2 | F | |||
| L | ||||
| H |
Notes. (a): Amplitude of power-law at wavenumber pix. (b): Spectral index of the power-law. (c): Mask used to clean the data. F, L and H stand for fiducial, light and heavy masks (see Sect. 2.2). (d): Median brightness at m of diffuse emission computed on the Field 1 - Field 2 difference maps.
4 Power spectra of PAH emission
In this section, we derive power spectra of the Pleiades PAH emission using JWST data. To obtain these spectra, we need to subtract the CIB residual. This is achieved through three distinct methods, each based on different hypotheses. In Sect. 4.1, we assume that the CIB residual is the same in our two fields, both close to and distant from Merope, and we use the latter for subtraction. In Sect. 4.2, we use the anisotropy of the Pleiades emission to subtract the isotropic CIB residual. Sect. 4.3 presents fits of the Field 1 spectra using two power laws to represent the PAH contribution and the CIB residual.
| Field | Mask | ||
| (a) | (b) | ||
| Field 1 | |||
| Cut 1 | F | ||
| L | |||
| H | |||
| Cut 2 | F | ||
| L | |||
| H | |||
| Field 2 | |||
| Cut 1 | F | ||
| L | |||
| H | |||
| Cut 2 | F | ||
| L | |||
| H | |||
Notes.(a): Parameters of the power-law fit to for in Eq. 8. (b): Masks used to compute power spectra. F, L and H stand for fiducial, light and heavy.
4.1 Power spectra from spatial difference
For each spectrum obtained for Cuts 1 and 2 of Field 1, we subtract the mean of the Field 2 power spectra for Cuts 1 and 2. The difference subtracts the CIB residual and a potential contribution from the Galactic background to the PAH spectra. This approach does not make use of the anisotropy fit, which restricts the range of bins that can be measured with relevant accuracy. The differences in spectra are presented in Fig. 7. The dashed lines represent power-law fits:
where the division by corrects for the signal attenuation by the JWST beam.
The fit parameters, the amplitude and the spectral index , are listed in Table 4. The table also includes the mean of the sky emission measured after source masking and OFF subtraction. Figure 7 shows the spectra, and Table 4 lists the fit results, for the three masks. The light and heavy masks are used to assess systematic uncertainties. The data points in Fig. 7 show a noticeable steepening of the spectra at high for the heavy mask. The difference in the specral index is slightly larger than the statistical error. This may be due to the lack of small-scale structure in the multiscale background used to fill in the masked areas.
In this work, we assume that on the spatial scales probed by the Pleiades JWST data, PAH emission traces the gas column density. For scales smaller than the line-of-sight depth of the emitting medium, the spectral index of the column density power spectrum is identical to that of the three-dimensional density field (Cho and Lazarian, 2003; Brunt et al., 2010). We therefore proceed under the assumption that the PAH power spectrum is a proxy for the gas density power spectrum.
The spectral index that we measure matches the reference spectral index for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). It is also consistent with the index inferred by Armstrong et al. (1995) for electron density fluctuations in the local interstellar medium by analyzing pulsar scintillation measurements on scales below mpc.
To compare with previous measurements of dust power spectra of the diffuse ISM, we refer to the spectra measured by Gibson (2007) in the Pleiades and by Miville-Deschênes et al. (2016) for an independent field at high Galactic latitude. These spectra, which were obtained by analyzing images of dust scattered light, are the closest observational match to our analysis in terms of angular resolution. For far-infrared dust emission, we also consider results derived Herschel observations by Miville-Deschênes et al. (2010) and Auclair et al. (2024) with angular resolutions of 30 and , respectively. The spectra reported in these four dust studies with spectral indices ranging from -2.7 to -2.95 are shallower than the two shown in Figure 7 .
4.2 Power spectra from anisotropy
In our second approach, we assume that the residual CIB is isotropic and obtain power spectra for each of the JWST images taking the difference . These differences are shown in Fig. 8. To highlight the difference in slope between the Field 1 and 2 images, the spectra are scaled by . The dashed-lines represent power-law fits to the spectra:
The fit parameters and for each of the images are listed in Table 5. We also list values for the light and heavy masks to assess systematic errors. The differences with the values for the fiducial mask are minor. It is satisfying to obtain values for the power law indices of the spectra, which are consistent with those obtained for the Field 1 cuts using the spatial difference (see Table 4).
4.3 Two components fit
We present a third approach where we fit the and spectra separately. In this fit, we assume that the CIB residual is isotropic and well approximated by a power law. The equations of the data model are as follows:
where the PAH power spectra are fitted with power laws that have distinct amplitudes and spectral indices.
We perform a Monte Carlo Markov chain analysis (MCMC) to sample the joint posterior distribution of the model parameters. We utilize the EMCEE python package as described by Foreman-Mackey et al. (2013) in accordance with their recommendations. We use a Gaussian likelihood. The priors force the three amplitudes , and to be positive, and the spectral indices and to be negative. In addition, we set a Gaussian prior in the spectral index with a dispersion of 0.2 (). This prior follows from the fit of for Field 2 images illustrated in Fig. 6. As this prior makes use of Field 2 data, we run the MCMC analysis only for the two Field 1 images.
The model parameters with their error bars are listed in Table 6. The table lists the median values of the posterior distribution and error bars computed from the 16 and 84 percentiles. As shown in Fig. 9 for Field 1 Cut 1, the model provides a good fit for both and . The amplitude of the CIB residual is consistent with that measured for parallel spectra of Field 2, within error bars (see Table 3). The values of also match the spectral indices derived in Sects. 4.1 and 4.2. It is satisfying to validate these values using a distinct approach in which the CIB residual is included in the data model with a power-law spectrum, rather than being subtracted by a difference. We note that the degree of anisotropy, measured by the ratios in Table 6, differs by a factor 2 between Cuts 1 and 2 .
Figure 10 shows the joint posterior distributions of and , derived from the MCMC fits of Field 1 Cuts 1 and 2. The difference between the two spectral indices depends on and, more generally, on the simplified model used to separate the PAH spectra from the CIB residual. In Figure 10, the median value of is slightly negative, but we consider the deviation from zero to be not statistically significant. We conclude that the two spectral indices are identical within a measurement error bar of 0.2 ().
As Lithwick and Goldreich (2001)777See Sect. 3 in this publication pointed out in the context of their interpretation of radio scintillation observations, the observational effects of the parallel direction are washed out due to variations in the local orientation of anisotropy along the line of sight and across the sky. In our observations, the PAH emission is so highly anisotropic that modes perpendicular to the local orientation of anisotropy dominate the Fourier spectra in all projected orientations onto the plane of the sky. This statement entails three key facts. First, the ratios in Table 6 are lower limits to the intrinsic anisotropy of the PAH emission. These limits depend on the dispersion of the local orientation of the anisotropy within the emitting medium. Second, the spectral index in Eq. 9 does not represent the spectral index of the energy cascade in the direction parallel to the magnetic field. Third, the finding that is consistent with 0 does not exclude the possibility that the spectral indices of the turbulent energy cascade for perpendicular and parallel modes differ.
| Image | ||||||
| Cut 1 | ||||||
| Cut 2 | ||||||
| Cuts 1 & 2 |
Notes. Parameters of the joint fits of and for Field 1 images Cuts 1 and 2, using data model in Eq. 9 with pix-1.
5 Relation to interstellar turbulence
In this section, we present three complementary perspectives that relate the JWST Pleiades observations to interstellar turbulence. Together, they provide a framework that we hope will motivate theoretical studies aimed at interpreting the data.
5.1 Dynamical interaction between the Pleiades stars and the surrounding ISM
The Pleiades reflection nebulosity displays an impressive array of striations, which are locally nearly parallel. The origin of these striations remains a long-standing question. The idea that they could result from the dynamical interaction between the Pleiades stars with the surrounding ISM is not new. It was first discussed by Arny (1977), and remains attractive. The dust and gas observations analyzed by Herbig and Simon (2001) and Ritchey et al. (2006) are direct evidence of a dynamical interaction. The change in dust polarization angles observed with the Planck data in Fig. 21 is further evidence of this.
Projection effects make it difficult to infer the distance at which the interaction with the stars occurs. Modeling of scattered light from dust indicates that the distance between matter and the brightest Pleiades stars ranges from a few tenths to 1 parsec. Specifically, it is a few tenths of a parsec near Merope (Gibson and Nordsieck, 2003b), a value that is consistent with our estimate of the radiation field intensity at the position of the Field 1 images (see Sect. C). For Merope’s luminosity, a value of G in the range 10–20 corresponds to a distance from the star between 0.3 and 0.5 pc.
Gordon and Arny (1984) discuss three ways in which the gas and stars could interact: radiation pressure on the dust, stellar winds, and thermal pressure of ionized gas. Each of these could possibly account for the lack of matter near the stars. The JWST observations show that the striations extend to scales three orders of magnitude smaller than the scale of the interaction. Given that the Reynolds number of interstellar turbulence is high (Elmegreen and Scalo, 2004), what we observe with the JWST is most likely the signature of a turbulent energy cascade driven by interaction, rather than the interaction itself. Energy transfer could, for example, be mediated by radiation pressure, as investigated by Callies et al. (2025).
5.2 CNM structure on the scales of the Pleiades JWST observations
The conventional framework for interpreting CNM observations (Kritsuk et al., 2018; Bellomi et al., 2020; Marchal and Miville-Deschênes, 2021) is based on numerical studies of the interplay between supersonic turbulence and thermal instability (Seifried et al., 2011; Kim et al., 2013; Hennebelle and Iffrig, 2014; Saury et al., 2014). The JWST data open a novel perspective by tracing the structure of the CNM on spatial scales where the turbulence turnover timescale is shorter than the CNM cooling timescale.
To support this statement, we compute the eddy turnover timescale for a power law scaling of the velocity dispersion:
where is the dispersion of turbulent velocity on the scale and is the scale of energy injection. We use the size of the Pleiades nebula for , and the dispersion of velocities among gas tracers reported by Ritchey et al. (2006) as an observational estimate of .
We assume that magnetic fields are dynamically important on the scale of the Pleiades nebula. We therefore use the canonical scaling of the velocity dispersion for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). For this index, the velocity dispersion is subsonic on the scales (mpc) statistically sampled by our JWST observations. The eddy turnover timescale is as follows:
We compare with the gas cooling time, which we computed for a reference pressure in the local ISM using the analytical expression of the cooling function in Jennings and Li (2021). For mpc, is at least two orders of magnitude smaller than the cooling timescale of the warm neutral medium. For mpc, is comparable to the cooling time of the CNM, and shorter for lower values of . It is also relevant to estimate the Field length . Thermal instability is suppressed by conduction on scales smaller than .
where is the thermal conduction coefficient, the gas density, the gas temperature, and the gas cooling rate (Jennings and Li, 2021). The numerical value is calculated for characteristic values of and in the CNM. We find that is larger than the scales probed in our data analysis. These calculations suggest that the JWST observations of the Pleiades probe spatial scales on which turbulence in the CNM is subsonic and is no longer coupled to the thermal instability. As such, the data offer a new angle for comparing observations and theoretical models.
5.3 MHD subsonic turbulence
In order to interpret our data in the context of turbulence, one would need to connect the observational findings of this work with theoretical models of MHD turbulence scaling laws and numerical simulations (see Schekochihin, 2022, for a recent review). This is obviously a difficult task and it is beyond the scope of our observational analysis to draw conclusions about specific models. Our aim is to guide future theoretical investigations. With this in mind, we present what we consider to be the four key observational results of our study.
The mean orientation of the map striations coincides with that of the mean magnetic field projected onto the plane of the sky, as determined by the Planck dust polarization observations.
The anisotropy power ratios relative to this mean orientation, 90 and 40 for Fields 1 and 2 (Table 6), are lower bounds to the local anisotropy, due to the dispersion in the orientation of the striations (Sect. 3.2). This dispersion – as determined by the widths of the angular profiles of the 2 dimensional power spectra – is significant (see Table 2).
The anisotropy of is observed to be scale independent within a small uncertainty set by the separation between the PAH emission and the CIB residual. We stress that for a large power anisotropy, we expect perpendicular modes to be dominant in all orientations projected onto the plane of the sky (Sect. 4.3)
The spectral indices of the PAH power spectra measured for Field 1 match the reference value -3.5 for MHD turbulence. However, we note that the values for Field 2 differ significantly from this reference.
We discuss two possible interpretations of the origin of the power anisotropy. This choice is illustrative and not exclusive of other models. (1) Striations are formed by fast magnetosonic waves traveling perpendicularly to the mean magnetic field. (2) They are fingerprints of the anisotropy of the turbulent energy cascade of subsonic MHD turbulence. In the first case, the anisotropy is global with respect to an average guide field, while the anisotropy of the turbulence energy cascade occurs with respect to the local orientation of the magnetic field.
The first interpretation was quantified by Beattie and Federrath (2020) using numerical simulations. More research involving dedicated simulations is required to verify whether this interpretation can reproduce the main observational results. The second is based on the Goldreich and Sridhar (1995) model of subsonic MHD turbulence. In this framework, compressible modes are sheared and passively mixed by Alfvén waves (Lithwick and Goldreich, 2001). This interpretation is attractive because it links the JWST observations to the anisotropic nature of the MHD turbulence energy cascade (Boldyrev, 2006; Schekochihin, 2022).
6 Conclusion
We have presented and analyzed NIRCam JWST observations of the Pleiades nebula in the F335M filter centered on the PAH emission feature at m. These JWST observations provide an opportunity to analyze the structure of the CNM in exceptional detail, reaching spatial scales where existing observations are too limited to enable a statistical analysis. The main results of our work are summarized as follows.
We build maps of diffuse PAH emission, separating it from sources, stars, and galaxies, as well as the extended spikes of bright stars. The stars were subtracted using an empirical template of the JWST PSF. Galaxies are identified and masked using multiresolution image filtering to interpolate the structure of diffuse emission.
We present a power spectrum analysis of the diffuse emission in four images, two of which are close to the Merope star and two of which are more distant. The spectra exhibit flattening at high wavenumbers, which we identify as a CIB residual. The spectral index of this component, -1.2, is similar to that measured for galaxy clustering in CIB observations at near- and far-infrared wavelengths.
The amplitude of PAH power spectra are highly anisotropic on all scales. We characterize them in three independent ways. (1) We subtract the CIB residual by taking the difference of the spectra computed for the JWST images near and far from Merope. (2) We differentiate anisotropic PAH emission from the isotropic CIB residual by modeling the 2D angular distribution of Fourier power. (3) We fit the power spectra in the directions parallel and perpendicular to the anisotropy orientation as the sum of two power laws, representing the contributions from PAHs and the CIB. The three methods yield consistent results.
The PAH power spectra are measured to a spatial resolution of 0.2 mpc (40 au). They are well fitted with single power laws, with no sign of a break that could be interpreted as a signature of a characteristic scale associated with turbulence dissipation. The power laws indices are -3.5 in the two fields near Merope and -3 in the more distant areas. The magnetic field orientation inferred from dust polarization data from Planck is aligned with the PAH anisotropy. The power anisotropy does not change with scale.
These findings are discussed in the context of interstellar turbulence, which may be driven by the dynamical interaction between the Pleiades stars and the surrounding ISM. The observations offer a new perspective for comparing observations and theoretical models because they reveal physical scales at which CNM turbulence is decoupled from the thermal instability. Progress in interpreting the data within this framework requires follow-up theoretical work including numerical simulations to link the PAH power spectrum to the three-dimensional, anisotropic structure of the magnetized ISM.
Acknowledgements.
GV, AAC and NDR thank the STScI for DDRF Summer funding to support their research. We thank the referee for a detailed report, which helped us to clarify the discussion of the results of the data analysis in relation to interstellar turbulence.References
- The Distance to the Pleiades According to Gaia DR2. Research Notes of the American Astronomical Society 2 (3), pp. 150. External Links: Document, ADS entry Cited by: §1, §3.1.
- Toward High-Precision Astrometry with WFPC2. I. Deriving an Accurate Point-Spread Function. PASP 112 (776), pp. 1360–1382. External Links: Document, astro-ph/0006325, ADS entry Cited by: §A.3.
- Empirical Models for the WFC3/IR PSF. Note: Instrument Science Report WFC3 2016-12, 42 pages External Links: ADS entry Cited by: §A.3.
- Electron Density Power Spectrum in the Local Interstellar Medium. ApJ 443, pp. 209. External Links: Document, ADS entry Cited by: §4.1.
- A model for the filamentary structure in the Pleiades reflection nebulosity.. ApJ 217, pp. 83–89. External Links: Document, ADS entry Cited by: §5.1.
- Separation of dust emission from the cosmic infrared background in Herschel observations with wavelet phase harmonics. A&A 681, pp. A1. External Links: Document, 2305.14419, ADS entry Cited by: §4.1.
- Filaments and striations: anisotropies in observed, supersonic, highly magnetized turbulent clouds. MNRAS 492 (1), pp. 668–685. External Links: Document, 1911.13090, ADS entry Cited by: §5.3.
- 3D chemical structure of diffuse turbulent ISM. I. Statistics of the HI-to-H2 transition. A&A 643, pp. A36. External Links: Document, 2009.05466, ADS entry Cited by: §5.2.
- The Per-Tau Shell: A Giant Star-forming Spherical Shell Revealed by 3D Dust Observations. ApJ 919 (1), pp. L5. External Links: Document, 2109.09763, ADS entry Cited by: Appendix C.
- Spectrum of Magnetohydrodynamic Turbulence. Phys. Rev. Lett. 96 (11), pp. 115002. External Links: Document, astro-ph/0511290, ADS entry Cited by: §5.3.
- Astropy/photutils: 1.8.0. Zenodo. External Links: Document, Link Cited by: §A.2.
- A method for reconstructing the variance of a 3D physical field from 2D observations: application to turbulence in the interstellar medium. MNRAS 403 (3), pp. 1507–1515. External Links: Document, 1001.1046, ADS entry Cited by: §4.1.
- Alfvénic Turbulence Beyond the Ambipolar Diffusion Scale. ApJ 805 (2), pp. 118. External Links: Document, 1412.3452, ADS entry Cited by: §1.
- Diagnosing Turbulence in the Neutral and Molecular Interstellar Medium of Galaxies. PASP 133 (1028), pp. 102001. External Links: Document, 2106.02239, ADS entry Cited by: §1.
- Dust-polarization Maps and Interstellar Turbulence. ApJ 839 (2), pp. 91. External Links: Document, 1608.08138, ADS entry Cited by: §1.
- Impact of the ambipolar diffusion in the structuration of the magnetic Rayleigh Taylor instability with oblique magnetic field. A&A 698, pp. A147. External Links: Document, 2506.19672, ADS entry Cited by: §5.1.
- Angular spectra of polarized Galactic foregrounds. New A Rev. 47 (11-12), pp. 1143–1149. External Links: Document, astro-ph/0306183, ADS entry Cited by: §4.1.
- The Anisotropy of Magnetohydrodynamic Alfvénic Turbulence. ApJ 539 (1), pp. 273–282. External Links: Document, astro-ph/0003403, ADS entry Cited by: §3.2.
- Neutral Hydrogen Structures Trace Dust Polarization Angle: Implications for Cosmic Microwave Background Foregrounds. Phys. Rev. Lett. 115 (24), pp. 241302. External Links: Document, 1508.07005, ADS entry Cited by: §1, §3.2.
- The global dust SED: tracing the nature and evolution of dust with DustEM. A&A 525, pp. A103. External Links: Document, 1010.2769, ADS entry Cited by: Appendix B, Appendix C, §2.2.
- The Atacama Cosmology Telescope: Galactic Dust Structure and the Cosmic PAH Background in Cross-correlation with WISE. ApJ 960 (2), pp. 96. External Links: Document, 2307.06352, ADS entry Cited by: §3.1.
- Infrared Emission from Interstellar Dust. IV. The Silicate-Graphite-PAH Model in the Post-Spitzer Era. ApJ 657 (2), pp. 810–837. External Links: Document, astro-ph/0608003, ADS entry Cited by: Appendix B, Appendix C, §2.2.
- Detection and Characterization of Cold Interstellar Dust and Polycyclic Aromatic Hydrocarbon Emission, from COBE Observations. ApJ 475 (2), pp. 565–579. External Links: Document, astro-ph/9610198, ADS entry Cited by: footnote 2.
- Interstellar Turbulence I: Observations and Processes. ARA&A 42 (1), pp. 211–273. External Links: Document, astro-ph/0404451, ADS entry Cited by: §5.1.
- Spitzer/IRAC and ISOCAM/CVF insights on the origin of the near to mid-IR Galactic diffuse emission. A&A 453 (3), pp. 969–978. External Links: Document, astro-ph/0604238, ADS entry Cited by: Appendix C, §1, footnote 2.
- emcee: The MCMC Hammer. PASP 125 (925), pp. 306. External Links: Document, 1202.3665, ADS entry Cited by: §4.3.
- A Calculation of Confusion Noise Due to Infrared Cirrus. AJ 103, pp. 1313. External Links: Document, ADS entry Cited by: §1.
- Multiscale Structure in Dust Reflection and Cold HI. In SINS - Small Ionized and Neutral Structures in the Diffuse Interstellar Medium, M. Haverkorn and W. M. Goss (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 365, pp. 59. External Links: Document, astro-ph/0703055, ADS entry Cited by: §1, §4.1.
- The Pleiades Reflection Nebula. I. Ultraviolet, Optical, and Far-Infrared Imaging Photometry. ApJ 589 (1), pp. 347–361. External Links: Document, ADS entry Cited by: §1.
- The Pleiades Reflection Nebula. II. Simple Model Constraints on Dust Properties and Scattering Geometry. ApJ 589 (1), pp. 362–377. External Links: Document, ADS entry Cited by: §1, §5.1.
- Toward a Theory of Interstellar Turbulence. II. Strong Alfvenic Turbulence. ApJ 438, pp. 763. External Links: Document, ADS entry Cited by: §5.3.
- An H I cloud in the Pleiades : a cloud-cluster collision ?. AJ 89, pp. 672–682. External Links: Document, ADS entry Cited by: §1, §5.1.
- Tiny-Scale Atomic Structure and the Cold Neutral Medium. ApJ 481 (1), pp. 193–204. External Links: Document, ADS entry Cited by: §1.
- Turbulent molecular clouds. A&A Rev. 20, pp. 55. External Links: Document, 1211.0637, ADS entry Cited by: §1.
- Simulations of magnetized multiphase galactic disc regulated by supernovae explosions. A&A 570, pp. A81. External Links: Document, 1405.7819, ADS entry Cited by: §5.2.
- Barnard’s Merope Nebula Revisited: New Observational Results. AJ 121 (6), pp. 3138–3148. External Links: Document, ADS entry Cited by: §1, §5.1.
- MASTER of the Cosmic Microwave Background Anisotropy Power Spectrum: A Fast Method for Statistical Analysis of Large and Complex Cosmic Microwave Background Data Sets. ApJ 567 (1), pp. 2–17. External Links: Document, astro-ph/0105302, ADS entry Cited by: §3.1.
- Damping of MHD turbulence in a partially ionized medium. MNRAS 527 (2), pp. 3945–3961. External Links: Document, 2306.10010, ADS entry Cited by: §1.
- Structure and Colors of Diffuse Emission in the Spitzer Galactic First Look Survey. ApJS 154 (1), pp. 281–285. External Links: Document, astro-ph/0406237, ADS entry Cited by: §1.
- Turbulence of a Conducting Fluid in a Strong Magnetic Field. Sov. Ast. 7, pp. 566. External Links: ADS entry Cited by: §4.1, §5.2.
- Thermal instability and multiphase gas in the simulated interstellar medium with conduction, viscosity, and magnetic fields. MNRAS 505 (4), pp. 5238–5252. External Links: Document, 2012.05252, ADS entry Cited by: §5.2, §5.2.
- Can the observed E/B ratio for dust galactic foreground be explained by sub-Alfvénic turbulence?. MNRAS 472 (1), pp. L10–L14. External Links: Document, 1707.06276, ADS entry Cited by: §1.
- New Measurements of the Cosmic Infrared Background Fluctuations in Deep Spitzer/IRAC Survey Data and Their Cosmological Implications. ApJ 753 (1), pp. 63. External Links: Document, 1201.5617, ADS entry Cited by: §3.3.
- Three-dimensional Hydrodynamic Simulations of Multiphase Galactic Disks with Star Formation Feedback. I. Regulation of Star Formation Rates. ApJ 776 (1), pp. 1. External Links: Document, 1308.3231, ADS entry Cited by: §5.2.
- Determination of inflationary observables by cosmic microwave background anisotropy experiments. Phys. Rev. D 52 (8), pp. 4307–4318. External Links: Document, astro-ph/9504054, ADS entry Cited by: §3.1.
- Inertial-Range Spectrum of Hydromagnetic Turbulence. The Physics of Fluids 8, pp. 1385. Cited by: §4.1, §5.2.
- Dust-Polarization Maps for Local Interstellar Turbulence. Phys. Rev. Lett. 121 (2), pp. 021104. External Links: Document, 1711.11108, ADS entry Cited by: §5.2.
- Heating of Interstellar Gas by Large Molecules or Small Grains. ApJ 335, pp. 769. External Links: Document, ADS entry Cited by: §1.
- Compressible Magnetohydrodynamic Turbulence in Interstellar Plasmas. ApJ 562 (1), pp. 279–296. External Links: Document, astro-ph/0106425, ADS entry Cited by: §4.3, §5.3.
- Thermal and Turbulent Properties of the Warm Neutral Medium in the Solar Neighborhood. ApJ 908 (2), pp. 186. External Links: Document, 2012.03160, ADS entry Cited by: §5.2.
- A Full-sky, High-resolution Atlas of Galactic 12 m Dust Emission with WISE. ApJ 781 (1), pp. 5. External Links: Document, 1312.0947, ADS entry Cited by: Table 7, Appendix B, Appendix C.
- A VLBI resolution of the Pleiades distance controversy. Science 345 (6200), pp. 1029–1032. External Links: Document, 1408.6544, ADS entry Cited by: §3.1.
- Probing interstellar turbulence in cirrus with deep optical imaging: no sign of energy dissipation at 0.01 pc scale. A&A 593, pp. A4. External Links: Document, 1605.08360, ADS entry Cited by: §1, §4.1.
- Statistical properties of dust far-infrared emission. A&A 469 (2), pp. 595–605. External Links: Document, 0704.2175, ADS entry Cited by: Appendix C, §1.
- Herschel-SPIRE observations of the Polaris flare: Structure of the diffuse interstellar medium at the sub-parsec scale. A&A 518, pp. L104. External Links: Document, 1005.2746, ADS entry Cited by: §1, §4.1.
- Physics and chemistry of interstellar polycyclic aromatic molecules.. A&A 164, pp. 159–178. External Links: ADS entry Cited by: §1.
- Updated point spread function simulations for JWST with WebbPSF. In Space Telescopes and Instrumentation 2014: Optical, Infrared, and Millimeter Wave, Jr. Oschmann, M. Clampin, G. G. Fazio, and H. A. MacEwen (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9143, pp. 91433X. External Links: Document, ADS entry Cited by: §A.3.
- Planck 2013 results. XI. All-sky model of thermal dust emission. A&A 571, pp. A11. External Links: Document, 1312.1300, ADS entry Cited by: Appendix B, Appendix C.
- Planck intermediate results. XLIV. Structure of the Galactic magnetic field from dust polarization maps of the southern Galactic cap. A&A 596, pp. A105. External Links: Document, 1604.01029, ADS entry Cited by: §3.2.
- Planck intermediate results. XLVIII. Disentangling Galactic dust emission and cosmic infrared background anisotropies. A&A 596, pp. A109. External Links: Document, 1605.09387, ADS entry Cited by: Appendix B.
- Planck intermediate results. XXXII. The relative orientation between the magnetic field and structures traced by interstellar dust. A&A 586, pp. A135. External Links: Document, 1409.6728, ADS entry Cited by: §1, §3.2.
- Planck 2018 results. XI. Polarized dust foregrounds. A&A 641, pp. A11. External Links: Document, 1801.04945, ADS entry Cited by: §1.
- Planck 2018 results. XII. Galactic astrophysics using polarized dust emission. A&A 641, pp. A12. External Links: Document, 1807.06212, ADS entry Cited by: Appendix B.
- Comparing the morphology of molecular clouds without supervision. A&A 696, pp. A217. External Links: Document, 2407.09832, ADS entry Cited by: §1.
- Performance of NIRCam on JWST in Flight. PASP 135 (1044), pp. 028001. External Links: Document, 2212.12069, ADS entry Cited by: §2.1.
- How Dark the Sky: The JWST Backgrounds. PASP 135 (1046), pp. 048002. External Links: Document, 2211.09890, ADS entry Cited by: §2.2.
- The Nature of Interstellar Gas toward the Pleiades Revealed in Absorption Lines. ApJ 649 (2), pp. 788–806. External Links: Document, astro-ph/0606644, ADS entry Cited by: §1, §5.1, §5.2.
- The structure of the thermally bistable and turbulent atomic gas in the local interstellar medium. A&A 567, pp. A16. External Links: Document, 1301.3446, ADS entry Cited by: §5.2.
- MHD turbulence: a biased review. Journal of Plasma Physics 88 (5), pp. 155880501. External Links: Document, 2010.00699, ADS entry Cited by: §5.3, §5.3.
- Forced turbulence in thermally bistable gas: a parameter study. A&A 526, pp. A14. External Links: Document, 1009.2871, ADS entry Cited by: §5.2.
- Atomic and Ionized Microstructures in the Diffuse Interstellar Medium. ARA&A 56, pp. 489–540. External Links: Document, 1810.00933, ADS entry Cited by: §1.
- A CO Survey of the Dark Nebulae in Perseus, Taurus, and Auriga. ApJS 63, pp. 645. External Links: Document, ADS entry Cited by: Appendix B.
- scikit-image: Image processing in Python. PeerJ 2, pp. e453. External Links: Document, 1407.6245, ADS entry Cited by: §A.4.
- HerMES: Cosmic Infrared Background Anisotropies and the Clustering of Dusty Star-forming Galaxies. ApJ 772 (1), pp. 77. External Links: Document, 1208.5049, ADS entry Cited by: §3.1.
- The Wide-field Infrared Survey Explorer (WISE): Mission Description and Initial On-orbit Performance. AJ 140 (6), pp. 1868–1881. External Links: Document, 1008.0031, ADS entry Cited by: §2.2.
Appendix A Obtaining images of the diffuse emission
This appendix details and illustrates the methods used to create maps of diffuse emission out of the JWST mosaics. Fig. 11 presents a zoomed-in view of the data to illustrate the diversity of emission sources we need to subtract or mask to obtain maps that can be used to statistically analyze the diffuse emission.
A.1 Removing diffraction spikes from off-field point sources
The bright star Merope in the vicinity of our field of view produces three diffraction spikes that are bright features across Field 1 (see bottom left panel of Fig. 1). The central one to the East of Merope aligns with the NIRCam detector arrays, and the other two are at . The Merope spike with the lowest declination extends to Field 2. We model and remove the three spikes from Merope in Field 1, the spike from Merope that extends to Field 2, and two fainter spikes from PQ Tau in Field 1. The first three spikes are clearly visible on the bottom right panel in Fig. 1, while the fourth is barely noticeable on the bottom left panel. The last two are too faint to be visible in the Figure. We illustrate our modeling with the eastern spike.
To model the oblique spikes, we rotate the JWST mosaic in array coordinates by plus or minus 60 degrees to apply the same procedure.
We cut a rectangular section of the observations surrounding the diffraction spike, which is 75 to 175 pixels wide on each side (see the first panel in Fig. 12). The width chosen depends on the intensity of the spike. In the following, columns refer to one-dimensional cuts within the cutout image, across the spike, and rows to one-dimensional cuts that are parallel to the spike. The steps involved in modeling and subtracting the spikes are illustrated in Fig. 12.
The first step is to estimate the background emission around the spike. To achieve this, we mask the inner 50 to 150 pixels where the spike dominates the emission. We also mask any value above from the median in each column (using sigma_clip from Astropy). The data along each column are fitted with a third-order polynomial function, which we use to interpolate and fill in the diffuse emission at the position of the masked pixels. We apply a median filter on the polynomial interpolation, with a width along the spike orientation of 100 pixels, to filter out irregularities due to compact sources. The resulting map is shown in the second panel of Fig. 12. We subtract this background from the initial cutout image, which yields a map of the spike. This map still contains point and compact sources, as well as some residual diffuse emission (see the third panel in Fig. 12).
The second step involves modeling the spike. We estimate the gradient along the diffraction spike, assuming that it can be modeled by one polynomial function that applies to each row of pixels. To obtain this polynomial function, we calculate the -clipped median of each column across the width of the spike: 150 pixels for the eastern spike of Merope, 100 pixels for the other spikes. This gives us the typical longitudinal gradient profile of the spike, which we fit with a 7th order polynomial function (see Fig. 13).
Finally, we normalize the first approximation of the spike we obtained earlier by dividing it, for each column across the spike, by the polynomial values of the longitudinal gradient. For each column, we then select the nearest 100 columns to compute the -clipped median of this selection for each row. This gives us the profiles across the spike, assuming it does not vary on scales much shorter than 100 pixels or 6.3″(see Fig. 14).
We note that the pattern varies smoothly along the spike. In particular, we see the peak of the profile shifts by about 2 pixels over 2000 pixels. A larger shift seems to occur at the faintest end of the spike, far from Merope, where the relative intensity of the spike is the lowest. However, even there, the shift is not greater than 7 pixels compared to the other end of the spike. Next, we rescale the normalized profiles with the polynomial gradient to obtain the spike model shown in the fourth panel of Fig. 12. Finally, we subtract the model from the initial map (see the fifth panel in Fig. 12). To assess the accuracy with which the spike is removed, we compared the residual background in the spike image, away from the central area, relative to the initial map. We find that, in the case of Merope’s straight spike, the residual is typically lower than 1% at more than 75 pixels (4.7 ″) from the center of the spike (see Fig. 15).
A limitation in our method is visible in Fig. 16. Closer to Merope, the pattern across the spike varies on scales smaller than 100 pixels (or 6.1″), leaving a residual trace near the edge of our observations. However, the alternative (i.e. using a smaller window than 100 pixels) would not properly filter out the sources “hidden” by the spikes. We compare these residuals with local dust emission by normalizing the profiles at the site of the spike to their median values (see Figure 17). In most cases, the residual of the spike is barely noticeable above the intrinsic variations of the diffuse emission, at a level of 5% of the median value of the local background. However, near Merope, the residual reaches more than 10% of that level but remains mostly random.
At this point, we thus have four mosaics, two for each field, with all large diffraction spikes removed. For each pair of mosaics, we remove the glitches by comparing both mosaics. Each pixel whose value varies by more than 0.5 MJy/sr is replaced by the lower value in the pair.
A.2 Multiscale background
Once the large spikes have been removed from the data, we model the diffuse emission. We used mosaics of Fields 1 and 2 from which the spikes have been removed. We compute their global median by clipping pixels above 2- and subtract the result from the images.
Next, we subtract the different scales of the diffuse emission by repeating three times the following steps: (1) we compute the standard deviation of the 2--clipped background subtracted image, (2) we mask the bright pixels whose values are more than 15- above the median of the entire image, (3) we perform a 2D median-filtering with a 4- clipping without the masked pixels, and (4) we subtract the median-filtered image. We use the Background2D class from photutils to perform these steps (Bradley et al. 2023). We apply the 2D median filter using rectangular boxes elongated along the direction of the detector arrays, which is close to the orientation of the striations. We use boxes with mean sizes of 100, 40 and 25 pixels. The ratio between the long and short sizes is 2 at all three scales. Figure 18 illustrates the composition of the background map as smaller scales are included, as well as the residual map obtained after background subtraction for a sub-image of Field 1. There is structure in the diffuse emission at scales smaller than 25 pixels (i.e. 1.5 arcsec). This contributes noticeably, albeit minimally, to the residuals, even in Field 1, where the diffuse emission is strongest.
A.3 Removing point sources within the field
After subtracting the multiscale background, we use the residual images to identify point sources and model the PSF. We identify point sources by recursively selecting the brightest pixel in the four combined mosaics. For each such pixel, we look for all other neighbor bright pixels and label the cluster as a source before looking for the next brightest pixel in the observations. We extract in this way the brightest 400 sources from our observations. We then visually inspect them all to separate point sources from extended sources, and end up with 175 point sources. We do not use publicly available source extraction algorithms as they tend to identify multiple sources within each JWST PSF, as those are highly structured for the brightest sources. Our method is more robust and still functional given the small number of sources we want to extract.
To build an empirical PSF (ePSF), we follow the method developed by Anderson and King (2000) and Anderson (2016), which is implemented as the EPSFBuilder class in photutils, adapting it to our specific needs. In particular, we want to clean the spikes of the point sources better than the cores because the former contaminate the diffuse emission more than the latter.
To combine the images of the selected point sources, we first center each image to subpixel accuracy using the theoretical PSF (tPSF) of WebbPSF (Perrin et al. 2014) on a 301 by 301 pixel ( by ). Each PSF is fitted in a squared annulus region, instead of its core, and we find that using a region of 45 by 45 pixels, with the 9 by 9 innermost pixels masked, gives the best results. The difference between the position of the brightest pixel and the computed centers of the PSFs is typically less than 0.4 pixel.
The centered PSFs are then stacked in an iterative way by our modified EPSFBuilder to generate the ePSF for our observations. Here, we use the same square annulus as before to align the individual PSFs with each other. The method also returns the individual best fit for each source in our collection, in their 301 by 301 original pixel grid, which we subtract from the observations, except for the six brightest point sources that have diffraction spikes extending beyond 20″ (see Fig. 19). There are four point sources in Field 1 and two in Field 2, for which the core is highly saturated. If we had removed their ePSF best fit, this would have led to significant residuals across the entire 301 by 301 pixel region. For these sources, we instead wait until we add back the multi-scale background, after which we remove their diffraction spikes in the same way we use for the stars off field (see Sect. A.1).
A.4 Masking the remaining sources
At this stage, we have removed all diffraction spikes from sources within and without our field of observations. We are left with masking the saturated core of bright point sources and compact sources.
To build the mask of the remaining sources, we first identify pixels that correspond to significant residuals after subtracting the ePSF fit. We compute the 2--clipped standard deviation for each 301 by 301 pixel image of a point source, keeping only values above a floor of 0.1 MJy/sr. Any pixel in the image whose value in the residual of the ePSF fit is greater than 10 times this standard deviation is flagged. Any pixel whose value in the ePSF fit is greater than 15 times this standard deviation is also flagged. These pixels represent less than 0.5% of all pixels in the four sub-mosaics and are the first included in a mask.
Next, we mask additional bright pixels. We start by creating a multiscale background in an identical manner to what was done in the previous subsection, but now starting from a cleaner map. We then select all pixels where the residual emission after subtraction of the multiscale background is larger than a brightness threshold (0.035 MJy/sr for the fiducial mask). We then identify small and large clusters of masked pixels: those with fewer than or more than nine connected masked pixels. We verify that the small clusters typically correspond to faint stars or isolated glitches. Large clusters correspond either to the core of galaxies, whose shapes and sizes are very diverse (see Figure 11), or to the few bright point sources for which the ePSF method was not applied. Since the brightness threshold leaves the edges of bright sources unmasked, we extend the large clusters by a buffer zone of fifteen additional pixels. To do this, we use the skimage.morphology package (van der Walt et al. 2014). Finally, we add back the small clusters of pixels to the mask.
Starting from mosaics containing large diffraction spikes, point sources, and extended sources, we have step by step removed the spikes from both the off-field and in-field sources and masked both the core of the point sources and the extended sources. This leaves us with mosaics containing only diffuse emission, which are used to perform the Fourier analysis presented in Sect. 3.
Appendix B Magnetized interstellar matter near the Pleiades stars
| Field | ||||||
| K | MJy sr-1 | MJy sr-1 | MJy sr-1 | MJy sr-1 | ||
| (a) | (b) | (c ) | (c) | (c) | (d) | |
| Field 1 | ||||||
| Cut 1 | ||||||
| Cut 2 | ||||||
| Field 2 | ||||||
| Cut 1 | ||||||
| Cut 2 | ||||||
| Backgrounde |
Notes. (a) Dust temperature from Planck data at 5’ resolution. (b) Dust opacity at GHz from Planck data at 5’ resolution. (c) Stokes , and at 353 GHz measured at the center of the JWST images on Planck maps smoothed to a beam. The values for Fields 1 and 2 are background subtracted. (d) Brightness at m measured at the center of the JWST images on the WISE all-sky map (Meisner and Finkbeiner 2014). The values for Fields 1 and 2 are background subtracted. (e) Median values over the background area. The error bars are the standard deviation of pixel values. Uncertainties on the median values are one order of magnitude smaller.
This appendix provides a broader context for the JWST observations. Planck and WISE data are used to characterize the magnetized interstellar matter associated with the Pleiades nebula.
We used the dust temperature and opacity maps derived from fitting the dust spectral energy distribution at far-infrared wavelengths to a modified black body (Planck Collaboration 2016b), and the dust polarization maps at 353 GHz (Planck Collaboration 2020b). We also used the full sky image produced by Meisner and Finkbeiner (2014) from the WISE data at m888Image downloaded from https://faun.rc.fas.harvard.edu/ameisner/wssa/healpix.html. The angular resolution of these maps is 5’. However, Stokes and are smoothed to 7’ in order to improve the signal-to-noise ratio of the dust polarization.
The Planck and WISE maps offer complementary insights into the interstellar medium. The dust temperature traces the localized heating of dust grains by the Pleiades stars, while the dust opacity indicates the gas column density. The conversion factor from to for translucent molecular gas is (Planck Collaboration 2014). The dust polarization angles are perpendicular to the magnetic field component in the plane of the sky. The WISE image traces the emission of PAHs, the brightness of which scales linearly with the product of the intensity of the radiation field and the gas column density (Draine and Li 2007; Compiègne et al. 2011). Measurements of these observational tracers computed at the center positions of the four JWST images and for the surrounding background are listed in Table 7. The background measurements correspond to the median value calculated for pixels within of Merope where is smaller than to exclude the Pleiades area.
Figure 20 shows the Planck maps of dust temperature and submillimeter opacity map at GHz overlaid on the m WISE image at 12″resolution999The image was downloaded from https://irsa.ipac.caltech.edu/applications/wise. At the distance of the Pleiades stars, the angular size of the image corresponds to 3 pc. The dust temperature exceeds the background value across the full extent of the mid-infrared emission from the nebula. The opacity contours delineate an interstellar cloud to the south of Merope. It corresponds to a small cloud identified by Ungerechts and Thaddeus (1987) in their survey with a peak CO(1-0) brightness of . The Merope star and the two JWST fields are located at the edge of this translucent cloud. The overlap between the contours of and indicates that the cloud is locally heated by Merope.
Figure 21 shows polarization angles, which have been rotated by to display the magnetic field orientation in the plane of the sky. Angles were computed from the Planck maps after the background contribution to the Stokes and maps was subtracted. Therefore, the figure illustrates the orientation of the magnetic field within the Pleiades nebula. It shows how the magnetic field changes orientation over the nebula, in particular over the cloud to the south of Merope. The black squares indicate the JWST image fields, and the red segments show the orientations of the striations (the angles listed in Table 2). Figure 21 shows a good match between the red and white segments. The polarization angles measured at the center position of the JWST images can be found in Table 2.
Appendix C Association with the cold neutral medium
In the paper introduction, we claim that the JWST Pleiades observations probe the structure of the CNM within a few parsecs from Merope. To support and discuss this statement, we use the Planck and WISE data presented in Appendix B. These data provide a global perspective on the JWST observations. The angular resolution of the Planck data corresponds to a scale of 0.2 pc at the distance of Merope.
Merope lies to the front of the large (150 pc in diameter) Per-Tau bubble that was identified by Bialy et al. (2021) using 3 dimensional stellar reddening data. The star is close to the outer surface of the shell that surrounds the bubble. On smaller scales, the dust emission data presented in Appendix B show that Merope and the two JWST fields are located at the edge of a translucent cloud, with a few parsecs diameter. Table 7 lists the values of the dust temperature and the dust opacity in the center positions of the four JWST images, as well as for the surrounding background. The values of the Planck Stokes parameters at 353 GHz, and the PAH emission at m, measured on the WISE map of the entire sky produced by Meisner and Finkbeiner (2014), are also listed in the table. We used these data to estimate the radiation field intensity and gas column density at the location of the JWST fields.
The radiation field intensity may be estimated from the dust temperature using the formula:
with a spectral index (Planck Collaboration 2014). The factor measures the radiation field intensity normalized to its value within the background ISM, away from the Pleiades stars. This formula yields values ranging from 6 to 9 at the position of the JWST fields. The opacity values listed in Table 7 include the Galactic background. To estimate the column density of the cloud associated with Merope, we need to subtract this contribution. This subtraction introduces a large uncertainty equal to the dispersion of opacity values measured over the background area, which is listed as the error bar in the background line of the table. Taking into account this uncertainty, we consider that the column density of the cloud at the positions of the JWST fields cannot be much larger than .
The m brightness allows os to crosscheck the estimates of and . In the solar neighborhood, the mean emission at m is for a column density of (Compiègne et al. 2011). For fixed PAH abundance and emission properties, this emission scales linearly with the product (Draine and Li 2007; Compiègne et al. 2011). Based on the m brightness values in Table 7, we estimate the product to be 16 and 7 for Fields 1 and 2, where is expressed in units of . These values suggest that the estimates of G and derived from dust temperature and opacity are somewhat underestimated. This could be due to beam smoothing and the superposition of cloud and background emissions, which were not considered when fitting the dust temperature and opacity. In any case, the estimates of and support our assertion that the JWST observations characterize emission from CNM within a few parsecs from Merope. The value of is 10 at a distance of 0.5 pc from Merope. The contribution to of warm gas with densities of a few tenths across this distance is negligible.
We also estimate the contribution of the Galactic background to the JWST power spectra. To achieve this, we use the scaling law established by Miville-Deschênes et al. (2007) when analyzing IRAS data. The fraction of power contributed by the Galactic background is estimated using the following formula:
where , the brightness in the center of a JWST image in the F335M filter, is obtained by multiplying the values and in Table 1. To obtain the background brightness in the F335M filter, we scaled the background brightness at m listed in Table 7 by the intensity ratio , as determined in the study of Flagey et al. (2006). We obtain values of % for the Field 1 images and between one and a few % for the Field 2 images. These low values support the link that the paper makes between the JWST power spectra and the CNM near the Pleiades stars.